# Sparse Covariance Neural Networks

### v1

@@ 2 Problem Formulation
$\begin{gathered}\|\mathbf{H}(\mathbf{\hat{C}})\!-\!\mathbf{H}(\mathbf{C})\|\!\leq\!\frac{P}{\sqrt{t}}\mathcal{O}\left(\!\!\sqrt{N}\!+\!\frac{\|\mathbf{C}\|\sqrt{\log(Nt)}}{\nu t}\!\right)\!\!=\!\beta\!~{}\textnormal{and}~{}\|\mathsf{\Phi}(\mathbf{x},\mathbf{C},\mathcal{H})\!-\!\mathsf{\Phi}(\mathbf{x},\mathbf{\hat{C}},\mathcal{H})\|\!\leq\!LF^{L\!-\!1}\!\beta\end{gathered}$
@@
This notion of stability follows an extensive line of research on stability of GNNs to generic graph perturbations (Gama et al., 2020b; Cerviño et al., 2022; Arghal et al., 2022; Parada-Mayorga & Ribeiro, 2021; Keriven et al., 2020; Ruiz et al., 2020), but it specifically considers covariance estimation errors, which is fundamental for VNNs. Indeed, since in practical applications the true covariance is not available and VNNs operate on an estimate of it, this notion of stability acts as a certificate of performance guarantee w.r.t. the ideal scenario when working with the true covariance matrix and characterizes the impact of data and model characteristics on the performance. More in detail, the output difference of VNN caused by the sample estimation error is bounded proportionally by the square root of the number of samples $t$. When $t$ increases, the bound approaches zero, ultimately, showing convergence to the true data distribution. This result is specific to covariance estimation errors, whereas generic GNN stability bounds do not decrease with the number of samples. This improves the stability w.r.t. PCA, which we provide in the following lemma.
Lemma 1.
Consider a true covariance matrix $\mathbf{C}$ and a sample covariance estimate $\mathbf{\hat{C}}$ with respective eigendecompositions $\mathbf{C}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}^{\mathsf{T}}$ and $\mathbf{\hat{C}}=\mathbf{\hat{V}}\mathbf{\hat{\Lambda}}\mathbf{\hat{V}}^{\mathsf{T}}$. Then, for any signal $\mathbf{x}$ with $\|\mathbf{x}\|\leq 1$, it holds with high probability that $\|\mathbf{V}^{\mathsf{T}}\mathbf{x}-\mathbf{\hat{V}}^{\mathsf{T}}\mathbf{x}\|\leq\mathcal{O}(t^{-1/2}(\min_{i,j\neq i}|\lambda_{i}-\lambda_{j}|)^{-1})$.
@@
Lemma 1 shows that PCA stability is inversely proportional to the smallest gap between covariance eigenvalues, leading to unstable behaviors when the eigenvalues are close. VNNs do not suffer from this as the covariance filter can exhibit a stable response to close eigenvalues at the expense of lower discriminability (Sihag et al., 2022). Because of this advantage in finite-data settings, VNNs have been shown effective in covariance-based learning tasks, both on static and temporal data (Sihag et al., 2024; 2023; Cavallo et al., 2024b; a). However, they have two major limitations. First, when the true covariance matrix is sparse, the finite-sample estimate contains spurious correlations (Baik et al., 2005; Paul, 2007), affecting the performance of the VNN significantly. Second, when the true covariance is dense, a VNN on the finite-sample estimate is limited by its high quadratic computational complexity in the data dimension, restricting its applicability to low-dimensional settings only.
@@
Paper objective. Our goal is to overcome the above limitations by sparsifying the sample covariance matrix for the VNN. While PCA-based processing sparsification has been studied (Bickel & Levina, 2008b; Deshpande & Montanari, 2016), its extension to the VNN is challenging because it is unclear: (i) how to perform sparsification depending on whether the true covariance matrix is sparse or dense; (ii) what are the effects of sparsification on VNN stability. When the true covariance matrix is sparse, we apply thresholding-based sparsification strategies for the sample covariance matrix, and prove that they result in stability improvements (Thms. 2-3). When the true covariance matrix is dense, we put forth a stochastic sparsification framework in a form akin to dropout (Def. 4), characterize its impact on the VNN stability (Thm. 4), and propose principled sparsification strategies based on these findings (Sec. 4.3). All proofs are collected in the appendix.
3 Sparse True Covariance
In applications involving brain and spectroscopic imaging, weather forecasting, and finance, the sample covariance matrix is notoriously prone to spurious correlations due to limited sample size (Bickel & Levina, 2008a; Jobson & Korkie, 1980; Ledoit & Wolf, 2003). Sparsity-based regularizers for the covariance estimation have been proposed as a backup (Bickel & Levina, 2008b). In this section, we consider the scenario where true convariance matrices are sparse, propose two sparsification strategies, hard and soft thresholding, and analyze their effects on the VNN embeddings. The choice of hard and soft thresholding instead of other regularized covariance estimations (Ledoit & Wolf, 2003; Bickel & Levina, 2008a; Bien & Tibshirani, 2011; Friedman et al., 2007) lies in the fact that thresholding yields a fast approach and allows theoretical tractability.
Instrumental tool for our analysis will be the frequency response of a covariance filter. By computing the graph/covariance Fourier transform of input and output signal of the covariance filter (cf. equation 1) we get $\mathbf{\tilde{u}}=\mathbf{\hat{V}}^{\mathsf{T}}\mathbf{u}=\mathbf{\hat{V}}^{\mathsf{T}}\sum_{k=0}^{K}h_{k}[\mathbf{\hat{V}}\mathbf{\hat{\Lambda}}\mathbf{\hat{V}}^{\mathsf{T}}]^{k}\mathbf{x}=\sum_{k=0}^{K}h_{k}\mathbf{\hat{\Lambda}}^{k}\mathbf{\hat{V}}^{\mathsf{T}}\mathbf{x}$, which, for the $i$-th entry, leads to $\mathbf{\tilde{u}}_{i}=\sum_{k=0}^{K}h_{k}\hat{\lambda}_{i}^{k}\mathbf{\tilde{x}}_{i}=h(\hat{\lambda}_{i})\mathbf{\tilde{x}}_{i}$. This indicates that the frequency response of the covariance filter is a polynomial $h(\lambda)=\sum_{k=0}^{K}h_{k}\lambda^{k}$ with frequency variable $\lambda$ specified on the eigenvalues $\hat{\lambda}_{i}$.
Definition 1 (Lipschitz covariance filter).
The covariance filter is Lipschitz with constant $P$ if, for every pair of eigenvalues $\lambda_{i},\lambda_{j}$, $i\neq j$, its frequency response satisfies: $|h(\lambda_{i})-h(\lambda_{j})|\leq P|\lambda_{i}-\lambda_{j}|$.
Lipschitz covariance filters control the variability of the frequency response through the constant $P$. Higher $P$ allows the frequency response to generate different outputs at close eigenvalues and improves the filter’s discriminability, but may degrade its stability as shown in the following.
3.1 Hard thresholding
@@
Hard thresholding removes the covariance entries below a given value and it has been shown to improve covariance estimation in high-dimensional low-data settings (Bickel & Levina, 2008b).
@@
Given the sample covariance matrix $\mathbf{\hat{C}}$ and a coefficient $\tau>0$, the hard thresholding function is $\eta(\mathbf{\hat{C}})_{ij}=\hat{c}_{ij}$ if $|\hat{c}_{ij}|\geq\tau/\sqrt{t}$, 0 otherwise.
The hard threshold is inversely dependent on the number of samples as $1/\sqrt{t}$. This follows the intuition that the non-thresholded estimator approaches the true covariance as $t$ increases, hence, the sparsification is less needed and disappears in the limit of $t\rightarrow\infty$. Hard thresholding provides a more reliable covariance estimate because it removes small spurious finite-sample errors and, thus, improves the performance of VNNs when the true covariance is sparse. We analyze the impact of hard thresholding on the VNN stability in the following theorem.
@@
Let the true covariance $\mathbf{C}$ belong to the sparse class $\mathcal{C}=\{\mathbf{C}:c_{ii}\leq M,\sum_{j=1}^{N}\mathbbm{1}[c_{ij}\neq 0]\leq c_{0},\forall i\}$ where $M>0$ is a constant, $\mathbbm{1}(\cdot)$ is the indicator function and $c_{0}$ is the maximum number of non-zero elements in each row of $\mathbf{C}$. Consider a hard-thresholded sample covariance matrix $\mathbf{\bar{C}}$ following Def. 2 with $\tau=M^{\prime}\sqrt{\log N}$ and $M^{\prime}$ large enough. With a high probability, it holds that
$\|\mathbf{H}(\mathbf{\bar{C}})\mathbf{x}-\mathbf{H}(\mathbf{C})\mathbf{x}\|\leq t^{-1/2}Pc_{0}\sqrt{N\log{N}}(1+\sqrt{2N})+\mathcal{O}\left(t^{-1}\right).$
Thm. 2 shows that hard-thresholded covariance filters –and consequently VNN from the results in Thm. 1– are stable to covariance estimation errors and that they converge to the respective filter and VNN operating with the true covariance as $1/\sqrt{t}$ in the number of data samples. Importantly, hard thresholding improves the stability with a tighter bound than dense VNNs (Sihag et al., 2022) and sparse PCA (Bickel & Levina, 2008b).
Comparison with dense VNN. Contrasting the results of Thm. 2 and Thm. 1, we see that both bounds decrease with the same order $\mathcal{O}(t^{-1/2})$. However, the hard-thresholded VNN bound depends linearly on the number of non-zero elements in a row $c_{0}$, whereas that of the dense VNN depends on the spectral norm $\|\mathbf{C}\|$. For sparse covariance, $c_{0}\ll\|\mathbf{C}\|$ and the hard thresholding improves stability.
Comparison with sparse PCA. To provide further insights on the stability of sparse covariance filters, we contrast it also with the stability of sparse PCA.
@@
Consider a true covariance matrix $\mathbf{C}$ and the thresholded sample covariance estimate $\mathbf{\bar{C}}$ with respective eigendecompositions $\mathbf{C}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}^{\mathsf{T}}$ and $\mathbf{\bar{C}}=\mathbf{\bar{V}}\mathbf{\bar{\Lambda}}\mathbf{\bar{V}}^{\mathsf{T}}$. Then, for any signal $\mathbf{x}$ with $\|\mathbf{x}\|\leq 1$, it holds with high probability that
$\|\mathbf{V}^{\mathsf{T}}\mathbf{x}-\mathbf{\bar{V}}^{\mathsf{T}}\mathbf{x}\|\leq t^{-1/2}(\min_{i}|\lambda_{i}-\lambda_{i+1}|)^{-1}c_{0}N\sqrt{2\log{N}}$ (2)
That is, the stability of sparse PCA is inversely proportional to the minimum eigenvalue gap. This term is not present in the stability of S-VNN, i.e., Thm. 2, as it is absorbed by the filter Lipschitz constant $P$: indeed, the covariance filter can exhibit a stable frequency behavior for close eigenvalues at the expense of lower discriminability, but the latter is compensated in the subsequent VNN layers that use cascades of filterbanks and non-linearities. Thus, hard-thresholded VNNs attain improved stability w.r.t. sparse PCA in sparse covariance setting.
3.2 Soft thresholding
When the data follows the spiked covariance model, such as in electrocardiogram or brain image data (Johnstone & Lu, 2009), soft thresholding has been studied as a better alternative to achieve more reliable covariance estimates (Deshpande & Montanari, 2016). In this model, data points follow $\mathbf{x}_{i}=\sum_{q=1}^{r}\sqrt{\beta_{q}}u_{q,i}\mathbf{v}_{q}+\mathbf{z}_{i}$, where $\mathbf{v}_{1},\dots,\mathbf{v}_{r}\in\mathbb{R}^{N}$ are orthonormal vectors with exactly $c_{0}$ non-zero entries of magnitudes lower-bounded by $\theta/\sqrt{c_{0}}$ for some constant $\theta>0$, $u_{q,i}\sim\mathcal{N}(0,1)$ and $\mathbf{z}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ are i.i.d., and $\beta_{q}\in\mathbb{R}_{+}$ is a measure of the signal-to-noise ratio. In this case, we sparsify the covariance estimate with soft thresholding.

### v2

@@ II-B Covariance Filters
$\begin{gathered}\|\mathbf{H}(\mathbf{\hat{C}})\!-\!\mathbf{H}(\mathbf{C})\|\!\leq\!\frac{Pk_{\textnormal{max}}}{t^{1/2-\epsilon}}\mathcal{O}\left(\!\!\sqrt{N}\!+\!\frac{\|\mathbf{C}\|\sqrt{\log(Nt)}}{k_{\textnormal{min}}t^{2\epsilon}}\!\right)\!\!=\!\beta\!\end{gathered}$
for $\|\cdot\|$ denoting the 22- or the spectral norm.
We make the constant $k_{\textnormal{max}}$ explicit in the bound of Thm.˜1, whereas in [11, Thm. 2] it is embedded in the assumption on the filter frequency response. We opt for this formulation for an easier comparison with our following analysis.
This notion of stability follows an extensive line of research on stability of GNNs to generic graph perturbations [12, 19, 20, 21, 22, 23], but it specifically considers covariance estimation errors, which is fundamental for covariance filters. Since in practical applications the true covariance is not available and covariance filters operate on an estimate, this notion of stability acts as a certificate of performance guarantee w.r.t. the ideal scenario when working with the true covariance matrix and characterizes the impact of data and model characteristics. More in detail, it identifies that the output difference of a covariance filter caused by the sample estimation error is bounded proportionally by the square root of the number of samples $t$. The bound increases with the Lipschitz constant $P$, since a higher $P$ allows the frequency response to generate different outputs at close eigenvalues, thus improving the filter’s discriminability, but leads to larger sensitivity to estimation errors, thus degrading its stability.
Covariance filters show increased stability w.r.t. that of PCA, which we investigate in the following lemma.
Lemma 1.
Consider the PCA with a true covariance matrix $\mathbf{C}$ and a sample covariance estimate $\mathbf{\hat{C}}$ of respective eigendecompositions $\mathbf{C}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}^{\mathsf{T}}$ and $\mathbf{\hat{C}}=\mathbf{\hat{V}}\mathbf{\hat{\Lambda}}\mathbf{\hat{V}}^{\mathsf{T}}$. Then, for any signal $\mathbf{x}$ with $\|\mathbf{x}\|\leq 1$, it holds with probability 1o(1)1-o(1) that $\|\mathbf{V}^{\mathsf{T}}\mathbf{x}-\mathbf{\hat{V}}^{\mathsf{T}}\mathbf{x}\|\leq\mathcal{O}(t^{-1/2}(\min_{j}|\lambda_{j}-\lambda_{j+1}|)^{-1})$.
We refer to [24] for details on the probability 1o(1)1-o(1), which gets closer to 1 as $t$ increases. Lemma˜1 shows that PCA stability is inversely proportional to the smallest gap between covariance eigenvalues, leading to unstable behaviors when the eigenvalues are close. Nevertheless, covariance filters do not suffer from this issue as they exhibit a stable response to close eigenvalues, which is modeled by their frequency response and Lipschitz constant $P$.
Figure 2: Covariance graph. Each feature is a node, the signal $\mathbf{x}$ becomes a graph signal and the edge weight is the covariance value among two features.
II-C Covariance Neural Networks
A VNN is a sequence of layers and the $l$-th layer assembles a filter bank of $F_{l-1}\times F_{l}$ covariance filters and a non-linear activation function $\sigma(\cdot)$, i.e.,
$\mathbf{u}^{l}_{f}=\sigma\left(\sum_{g=1}^{F_{l-1}}\mathbf{H}^{l}_{fg}(\mathbf{\hat{C}})\mathbf{u}^{l-1}_{g}\right)~f=1,\ldots,F_{l},~l=1,\ldots,L$ (4)
@@ II-C Covariance Neural Networks
Theorem 2 ([11]).
Consider a VNN $\mathsf{\Phi}(\mathbf{x},\mathbf{C},\mathcal{H})$ of $L$ layers with $F_{l}=F$ for $l=1,\ldots,L$ and nonlinearities $\sigma(\cdot)$ such that $|\sigma(a)-\sigma(b)|\leq|a-b|$. Let $\beta$ be the largest stability bound of all covariance filters in the VNN as per Thm.˜1. Consider a generic data sample $\mathbf{x}$ with covariance $\mathbf{C}$ and $\|\mathbf{x}\|\leq 1$ w.l.o.g.. Then, with probability 1o(1)1-o(1) it holds that
$\begin{gathered}\|\mathsf{\Phi}(\mathbf{x},\mathbf{C},\mathcal{H})\!-\!\mathsf{\Phi}(\mathbf{x},\mathbf{\hat{C}},\mathcal{H})\|\!\leq\!LF^{L\!-\!1}\!\beta.\end{gathered}$
@@
With stability in finite-data settings, VNNs have been shown effective in covariance-based learning tasks, both on static and temporal data [13, 14, 15, 16]. However, they have two major limitations. First, finite-sample covariance estimates often contain spurious correlations [10, 9], affecting the performance of the VNN significantly. Second, a VNN on the finite-sample estimate is limited by its quadratic computational complexity in the data dimension, restricting its applicability to low-dimensional settings only.
II-D Contributions
@@
To overcome the above limitations, we propose Sparse VNNs (S-VNNs), which are VNNs operating on sparsified sample covariance matrices. While PCA-based processing sparsification has been studied [25, 18], its extension to the VNN is challenging because it is unclear: (i) how to perform sparsification appropriately based on the condition of true covariance matrix (i.e., whether the true covariance matrix is sparse or dense); (ii) what is the impact of sparsification on the VNN stability, which calls for an analysis beyond the small perturbation studied in [11] since the sparsification perturbation might be large. We address these aspects by studying different sparsification strategies for S-VNNs in accordance with different conditions of true covariances, and analyze their impact on stability. More in detail, when the true covariance matrix is sparse, we study thresholding-based sparsification strategies for the sample covariance matrix, and show that they result in tighter stability bounds, stronger robustness, and improved computational complexity (Thms. 3-4). When the true covariance matrix is dense, we put forth a stochastic sparsification framework in a form akin to dropout (Def. 4), characterize its impact on the VNN stability (Thm. 5), and propose principled sparsification strategies based on these findings (Sec. IV-B). All proofs are collected in the supplement.
III Sparse True Covariance
In applications involving brain and spectroscopic imaging, weather forecasting and financial business, the correlations between data points are generally sparse (e.g., only some brain regions activate simultaneously, only some stock prices are affected by similar factors, etc.), leading to sparse true covariance matrices [26, 27]. However, the sample covariance matrix is notoriously prone to spurious correlations due to limited sample size, resulting in an inappropriate dense estimates [28, 29, 30]. In this case, we leverage thresholding-based sparsification strategies, i.e., hard and soft thresholding, on the sample covariance matrix for S-VNNs and study their effects. Compared to other regularized covariance estimations [30, 28, 31, 32], the hard and soft thresholding provide computational efficiency and theoretical tractability.
III-A S-VNNs with Hard Thresholding
@@
Hard thresholding removes the covariance entries below a pre-defined value and it has been shown to improve covariance estimation in high-dimensional low-data settings [25].
@@
Given the sample covariance matrix $\mathbf{\hat{C}}$ and a coefficient $\tau>0$, the hard thresholding function is $\eta(\mathbf{\hat{C}})_{ij}=\hat{c}_{ij}$ if $|\hat{c}_{ij}|\geq\tau/\sqrt{t}$, and 0 otherwise.
Following [25], hard thresholding is inversely dependent on the number of samples as $\tau/\sqrt{t}$. This follows the intuition that the non-thresholded estimator approaches the true covariance as $t$ increases, hence, the sparsification is less needed and disappears in the limit of $t\rightarrow\infty$. Hard thresholding provides a more reliable covariance estimate because it removes small spurious finite-sample errors and, thus, improves the performance of VNNs when the true covariance is sparse. We then characterize the impact of hard thresholding on VNNs with the stability analysis in the following theorem.
@@
Let the true covariance $\mathbf{C}$ belong to the sparse class $\mathcal{C}=\{\mathbf{C}:c_{ii}\leq M,\sum_{j=1}^{N}\mathbbm{1}[c_{ij}\neq 0]\leq c_{0},\forall i\}$ where $M>0$ is a constant, $\mathbbm{1}(\cdot)$ is the indicator function and $c_{0}$ is the maximum number of non-zero elements in each row of $\mathbf{C}$. Let the eigenvalues of the true covariance $\{\lambda_{i}\}_{i=1}^{N}$ be all distinct and strictly positive. Consider a hard-thresholded sample covariance matrix $\mathbf{\bar{C}}$ following Def. 2 with $\tau=M^{\prime}\sqrt{\log N}$ and $M^{\prime}$ large enough and a covariance filter $\mathbf{H}(\cdot)$ that is Lipschitz with constant $P$. With probability 1o(1)1-o(1), for a large enough constant $C$, it holds that
$\|\mathbf{H}(\mathbf{\bar{C}})\mathbf{x}-\mathbf{H}(\mathbf{C})\mathbf{x}\|\leq$
$t^{-1/2}Pc_{0}\sqrt{N\log{N}}(1+\sqrt{N})+\mathcal{O}\left(t^{-1}\right).$
Thm.˜3 shows that hard-thresholded sparse covariance filters –and consequently S-VNN from Thm. 2– are stable to perturbations on both the finite-data covariance estimation error and the thresholding and that they converge to the respective filter and VNN operating over the true covariance at a rate of $1/\sqrt{t}$, as the number of data samples $t$ increases. The threshold parameter $\tau$ is set to a value that allows for the recovery of the true covariance sparsity pattern with the desired rate, therefore it does not appear in the bound (see [25] for details). Importantly, hard thresholding provides a tighter bound than nominal VNNs [11] and hard-thresholded sparse PCA [25], as we elaborate next.
Comparison with nominal VNN. From Thm. 3 and Thm. 2, we see that both bounds decrease with the same order $\mathcal{O}(t^{-1/2})$ (for $\epsilon\rightarrow 0$ in Thm. 2). However, the stability bound of hard-thresholded S-VNN depends linearly on the number of non-zero elements in a row $c_{0}$, whereas that of the VNN depends on the spectral norm $\|\mathbf{C}\|$. For sparse covariance, $c_{0}\ll\|\mathbf{C}\|$ and the stability bound is tighter.
Comparison with sparse PCA. To provide detailed insights on the stability comparison between sparse covariance filters and sparse PCA, we first analyze the stability of sparse PCA with hard thresholding in the following proposition.
@@
Consider a true covariance matrix $\mathbf{C}$ and the thresholded sample estimate $\mathbf{\bar{C}}$ with respective eigendecompositions $\mathbf{C}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}^{\mathsf{T}}$ and $\mathbf{\bar{C}}=\mathbf{\bar{V}}\mathbf{\bar{\Lambda}}\mathbf{\bar{V}}^{\mathsf{T}}$. Then, for any signal $\mathbf{x}$ with $\|\mathbf{x}\|\leq 1$, it holds with probability 1o(1)1-o(1) that
$\|\mathbf{V}^{\mathsf{T}}\mathbf{x}-\mathbf{\bar{V}}^{\mathsf{T}}\mathbf{x}\|\leq t^{-1/2}(\min_{j}|\lambda_{j}-\lambda_{j+1}|)^{-1}c_{0}N\sqrt{2\log{N}}.$
The stability of sparse PCA is inversely proportional to the minimum eigenvalue gap. This term is not present in the stability of S-VNN, i.e., Thm. 3, as it is absorbed by the filter Lipschitz constant $P$. This indicates that the sparse covariance filter can exhibit a stable spectral behavior for close eigenvalues at the expense of lower discriminability (due to the Lipschitz property [Def. 8]), but the latter is compensated in the subsequent S-VNN layers that use cascades of filterbanks and non-linearities to increase the model expressivity [33, 34]. Thus, hard-thresholded S-VNNs attain improved stability compared with sparse PCA in the sparse covariance setting.
III-B S-VNNs with Soft Thresholding
When data follows a spiked covariance model, such as electrocardiogram or brain image data [35], soft thresholding has been studied to achieve more reliable covariance estimates [18]. Specifically, data points in this model follow $\mathbf{x}_{i}=\sum_{q=1}^{r}\sqrt{\beta_{q}}u_{q,i}\mathbf{v}_{q}+\mathbf{z}_{i}$, where $\mathbf{v}_{1},\dots,\mathbf{v}_{r}\in\mathbb{R}^{N}$ are orthonormal vectors with $c_{0}$ non-zero entries whose magnitudes are lower-bounded by $\theta/\sqrt{c_{0}}$ for some constant $\theta>0$, $u_{q,i}\sim\mathcal{N}(0,1)$ and $\mathbf{z}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ are independent and follow normal distributions, and $\beta_{q}\in\mathbb{R}_{+}$ is a measure of the signal-to-noise ratio. In this case, we sparsify the covariance estimate with soft thresholding.